wiki:UC20210504

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 05/04/2021


User feedback and comments from institutional representatives

  • BYU - had some recent ARCSAT issues, everything fine with 3.5m
  • FGCU - nothing to report
  • Wyoming - no report
  • Oklahoma - nothing to report
  • Georgia State - nothing to report
  • NAPG - nothing to report
  • Seattle - nothing to report
  • UVa - nothing to report
  • JHU - nothing to report
  • Colorado - a recent undergrad course got good data with DIS
  • NMSU - nothing to report
  • Washington - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 4/01/2021 - 4/30/2021

0) Overview

APO transitioned to a 50% staffing limit on 4/21/21 when Otero county moved from the Yellow to Green COVID tier. The State of New Mexico evaluates the tiered positions every two weeks on Wednesdays for each county in NM. Otero County moved to the Green on 4/21; the next update will be on 5/5/21. Currently, Otero County has a 2.54% positivity rate and 7.2 cases per 100k rate. For comparison, Dona Ana, which is also in the Yellow tier, has a 4.76% positivity rate and 12.3 cases per 100k rate.

The data and numbers are available here: https://cv.nmhealth.org/public-health-orders-and-executive-orders/red-to-green/

We are still restricting all visitors to the site. Following the State of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors.

We are preparing to start having visitors to the site, following the State of New Mexico and NMSU guidelines. Weather for the month of April has been warming, with occasional wintery weather.

1) Telescope

The telescope is operational with no major issues this last month. 

2) Instruments

DIS: Scattered light: Red camera is starting to change but is still considered good. The blue camera is starting to change, and is also still considered good. As temperatures warm up, the red chip temperature needs to be monitored, and darks may need to be taken during cals.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. Instrument servicing starts on 5/4/21.

NICFPS: powered on, cold and operational. 

ARCTIC: powered on, cold, and operational. Issue with dome flats is being debugged and solutions are being worked on. The “fence” baffle on the mirror cover petal design is on site from UVa machine shop. Testing will take place hopefully in the coming weeks. The new rotation redesign for the diffuser was worked on; some issues were identified and parts are being shipped back to UW for modification.

3) ARCSAT

ARCSAT is operational and has been used for observations this past month. Off-axis guiding is available and is recommended for exposures greater than 10 minutes. Shelyak eShel commissioning is ongoing currently, and is also scheduled for May 17-30.


Additional telescope and instrument discussion

We will have a new 3.5m Observing Specialist coming on board soon, so users should be extra nice to them during training.

Users may be starting to wonder about if/when they can have class visits to APO this fall, so the sooner we know of people’s intentions or desires on that front, the better. We are in the process of evaluating what we can and cannot do this fall with respect to visitors, recognizing that this is a fluid situation as the New Mexico public health orders are still evolving. We will continue to keep the UC informed as needed. In the mean time, if users anticipate bringing groups of students to APO any time in Q3 (or even Q4), please let Russet, Ben, and Nancy know so they can fold this into their advance planning efforts.


Discussion of summer shutdown status

Realuminization of the 3.5m primary mirror was postponed last summer due to COVID. It is not yet clear whether it will be possible during the summer 2021 shutdown; if not, it will either need to happen during a non-summer time, or it will need to wait until summer 2022. We will have more information concerning the shutdown schedule soon. [Note added in proof: the summer shutdown dates are scheduled, and will be July 19-29. The last night of science operations prior to shutdown is July 18, and the return to science is planned for July 30. The return date is contingent upon the completion of sufficient on-sky engineering prior to that, so that date is subject to change depending on weather conditions. The primary mirror will NOT be realuminized this summer.]


KOSMOS Update

Work is continuing on detector wiring.


ARCTIC Diffuser Rotator Update

The part was sent back to UW for some additional engineering.


Q2 3.5m scheduling

There is still some open time available in the Q2 schedule. Users are reminded to follow the standard procedure of emailing requests for this time to Ben, Russet, Nancy, and their institutional scheduler.


Q2 ARCSAT scheduling

There is no open time available in Q2.


Q3 3.5m scheduling

The institutional allocations will be sent out later this week, with proposal deadlines of May 27.


Reminder about non-student training sessions

UC reps are asked to remind their users that requests for training during an upcoming quarter MUST be submitted no later than the time of the proposal submissions for that quarter. Such requests should be sent by email to Ben, Russet, Nancy, and Candace.


3.5m Science Symposium

SAVE THE DATE: The virtual APO science symposium will be held during the week of 26-30 July. This symposium is intended to showcase results from observations made with the 3.5m and ARCSAT telescopes. Students are particularly encouraged to present their research. The specific days and times are still being decided upon (it will likely be 2-3 days, ~ 4-5 hours each day), and more details will be announced in the next couple of weeks. Stay tuned!


ACTION ITEMS

Open action items from previous meetings:

  • Nancy - Develop a survey to get feedback on the virtual training sessions, distribute to all recent trainees. Status: CLOSED. Received 17 responses so far, from people at OU, CU, JHU, and BYU.
  • UC reps - notify DIS users of the potential for increased background and steps to check/remediate the issue. Status: CLOSED.
  • UC reps - notify ARCSAT users of SurveyCam status. Status: CLOSED.
  • Nancy - finalize date(s) for virtual science symposium so we can announce it no later than next month. Status: CLOSED (see above).

New action items from this meeting:

  • UC reps - remind users of deadline for requesting non-student training session in the upcoming quarter.

All Other Business

None.


Next meeting

The next meeting will be on Tuesday June 1 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.